LON01 – POP-C002.017 – 06/04/2016 – 10:00 – 12:00 – *Emergency Work* *COMPLETE*

Further to our NOC notice posted on the 31/03/2016 in respect to power at one of our POPs “C002.017 – Goswell Road”, Our network monitoring has alerted us to another loss of power on the secondary feed at this cab.

Structured engineers are attending site tomorrow morning within the above maintenance window to install equipment that will allow us to isolate the faulty hardware without further risk to services provided via this cabinet going forward. The work will involve the replacement of various power distribution hardware. At this time the POP is operating on its redundant power feed and all services have been re-routed where possible. Ethernet services provided by this cab are considered “at risk” until power has been fully restored. Transit services will automatically re-route in the event of a failure.

Due to the nature of the works, engineers will be working within a live cabinet. No issues are expected and extreme care will be taken while the works are under-taken.

Further updates will be provided in the morning.

UPDATE01 – 10:10
Engineers have started work.

UPDATE02 – 10:55
The new power distribution hardware has been installed and engineers will begin to power up the affected hardware 1 device at a time. Level3 are on site with us in the evn of another problem.

UPDATE03 – 11:10
All hardware has been powered up and the faulty device found (all be it with a bang and fire) Unfortunately the failed hardware is the redundant power supply on the network core within this rack. Redundant hardware of this size is not kept on site and we are currently in the process of sourcing another unit. Further updates to follow.

UPDATE04 – 12:02
Further test have been done and confirmed the PSU unit has failed. Engineers have removed a power supply from another unit from our 4th floor suite and installed it within the 2nd floor POP to confirm it is un-damaged by the recent events. Further updates to follow.

UPDATE05 – 12:45
Engineers have ordered a same day replacement from one of our suppliers. Engineers are going to remain on site to fit and commission the new hardware on arrival

UPDATE06 – 15:25
Engineers remain on site awaiting hardware. ETA was 15:30, however this has been pushed back due to a crash on the A3.

UPDATE07 – 15:25
Enginners remain on site having little fun. We have been advised the part is now in London and will be with us by 17:30

UPDATE08 – 17.35
The replacement part has arrived on site

UPDATE09 – 17.50
Despite our best efforts, our supplier has shipped the wrong part! Discussions had with them have concluded with no further delivery options today. New hardware has been sourced and is being made avaliable to site for a timed delivery. Engineers are attending again in the morning to swap out the (hopefully) correct new part. We do applogise for the delay in getting this resolved, however want to remind customers who route via this device that it is still operating as expected on its redundant supply.

UPDATE10 – 09:30 – 07/04/2016
Engineers have retured to site and are awaiting delivery of the new PSU.

UPDATE11 – 09:46 – 07/04/2016
Delivery update to advise the hardware will be on site before 11am.

UPDATE12 – 10:33 – 07/04/2016
Hardware has arrived on site and engineers have confirmed it is the correct unit this time.

UPDATE13 – 10:47 – 07/04/2016
Engineers have installed the new power supply and confirmed its operation within the core. A series of load tests have been conducted with normal operation observed.

UPDATE14 – 11:00 – 07/04/2016
We are happy the new power supply is operating as expected, however will continue to monitor its operation for the next few hours. The site is no longer classed as “at risk” and this issue will now be closed off.
We apologise one again for the delay this has taken to resolve and will be reviewing our internal procedures on hardware spares of this nature at Goswell Road.

LON01 – EasyXDSL – 24/03/2016 – 14:30 – *Emergency Maintenance*

We have just been made aware that our carrier in conjunction with BT will be conducting some emergency maintenance on several FTTC gateways within the South East area. This will cause connections to drop, however should reconnect almost instantly.

We apologize for the short notice given for these works and will advise once complete.

LON01 – EasyHTTP – 02/03/2016 – 17:30 – *Maintenance* *Complete*

Our network monitoring has alerted us to a memory problem with “Server01” within our EasyHTTP platform. The server is on-line and processing requests, however is becoming more unresponsive to our monitoring systems as time progresses.

To avoid a complete failure of the system, the decision has been made to power cycle the server at 17:30

SMTP and WEB services on this server will be unavailable for the duration. We will review the system for stability once the reboot is complete and take further action where required.

We apologise for any inconvenience this may cause.

LON01 – EasyIPT – 29/02/2016 – 21:30 till 22:00 *Emergency Maintenance* *Complete*

Following on from reports today of SIP channel limits being reached for both inbound and outbound calls affecting some of our managed PBX systems, work has been undertaken on our core softswitch to try and identify the root cause of this issue.

Changes have been made to the platform however a emergency reboot is required. Due to the size of the platform this reboot will take up-to 15 minutes to complete. During this time inbound and outbound calls will be limited.

We will advise once service is restored and calls are routing again.

Apologies for any inconvenience this may cause.

UPDATE01 – 21:40 softswitch reboot is complete, we are monitoring traffic flow

UPDATE02 – 22:55 A random check on managed PBX systems are showing registration, however this is not reflected on our softswitch, traffic is flowing however.

UPDATE03 – 22:10 Systems continue to show inaccurate data about the status of some registrations. We suspect stuck SIP sessions on some managed PBXs as this is affecting a wide range of PBX software versions.

UPDATE04 – 21:40 All affected managed PBXs are showing as on-line correctly.

LON01 – 11/11/2015 – 23:00 till 03:00 – EasyXDSL – Emergency Maintenance *COMPLETE*

We have been advised this afternoon by one of our interconnect providers that that they will be conducting a software upgrade on there core Junos router at Telehouse North between 23:00 and 03:00 tonight. This upgrade work will affect one of our fibre waves we use for DSL termination back in to our network at Goswell Road. DSL traffic traversing over this link will drop during the maintenance window, however should re-establish over an alternative link via Telehouse East which is unaffected.

The nature of these works include a reload of the Junos hardware at Telehouse North by our wave supplier and in theory should be the only disruption we see. Unfortunately sort disruptions like this will cause multiple DSL PPP session re-terminations as they flip back and forth between Telehouse North & Telehouse East as our network is designed to use the former where possible. Due this concern and avoiding any other unforeseen issues that will cause session drops during the maintenance windows, we will be manually shutting down the Telehouse North link at 22:50 tonight with the view of bringing it backup at 07:00 tomorrow morning. This will limit the session drops to these times and provide a stable service throughout.

We apologise for any inconvenience this may cause.

UPDATE 01 – 18:45

We have reviewed the original manual switch over time and concluded it does not allow any contingency should this be required. We have therefor updated this to 22:00
UPDATE 02 – 07:10 – 12/11/2015

This work is complete.

LON01 – EasyIPT – 13/07/2014 – 20:00 till 21:00 *Planned Maintenance* *COMPLETE*

We are aware some users have been seeing stuck SIP sessions today, whereby “BYE” messages have not been completely clearing down resulting in accounts hitting a concurrent call limit. A work around has been put in place for this until this evening when we can reload the core device to drop any sessions still in this state.

UPDATE01 – 20:05
This work has started and the server has been requested to reload.

UPDATE02 – 20:15
The server has reloaded and is accepting SIP requests, however is not processing any call routing.

UPDATE03 – 20:35
We have discovered that the Kamailio Proxy service has failed to start which is why calls are not being processed. We are working to restore the service.

UPDATE04 – 20:40
The service has been restored and test calls are now routing OK.

LON01 – 26/06/2015 – 13:00 – CORE01.LON01 *EMERGENCY WORKS* COMPLETE*

Our network monitoring has alerted us to a intermittent fan failure on PSU 1 within CORE01.LON01. This device has redundant PSUs so no impact is expected.

Engineers will attend this afternoon to swap out the failing power supply. Again no impact is expected to be seen, however the device will be marked as “at risk” for the duration enginners are working on it.

Updates will be provided once available.

UPDATE01 – 14:00

This work has been completed without any issue.

Internal Systems *RESOLVED*

We are currently experiencing issues with some internal systems that is preventing various customer service / support operations. Engineers are on route to the data centre, however local network rail cancellations and disruptions are delaying engineers from attending.

Please also note the above is effectiving email delivery so we are unable to process support tickets. We advise calling for any urgent issues.

Apologies for any inconvenience this may cause. Once we know more we will update this notice.

UPDATE01 – 12.00
Engineers have arrived on site.

UPDATE 02 – 12.30
Internal systems have been restored

EasyHTTP – 24/04/2015 – 20:00 – 20:30 – Emergency Maintenance *COMPLETE*

We are aware of a permission issue on the Plesk control panel for EasyHTTP which is generating “Access Denied” errors when changing various account settings.

A patch has been released and installed, however this does require a full reload of the server which will take place tonight at 20:00. Due to the size of the system this reload may take up to 20 minutes to complete. During this time all services will be unavailable.

LON01 – 05-04-2015 – 20:00 – 21:00 – Emergency Network Maintenance *Complete*

We have discovered that one of our internal BGP filters is incorrectly configured. We are currently reviewing all filters to ensure this is only limited to 1 instance. Once the review is complete we will be taking steps to correct the configuration issues found. This will result in routing table changes and disruption to various services.

We apologies for any inconvenience this may cause.

UPDATE 01 – 20:20
After reviewing:

core01.lon01.structuredcommunications.co.uk
core02.lon01.structuredcommunications.co.uk
access01.lon01.structuredcommunications.co.uk

We have discovered only 1 miss-configured BGP filter on core01.lon01.structuredcommunications.co.uk. The BGP filter was only client facing and not carrier as originally suspected. No changes have been made to our carrier peering and no loss of service was seen.